TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-11 23:36:00
subject: 2\24 Pt-1 ISS On Orbit Status 24-02-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
2\24 ISS On-Orbit Status 24-02-2003
Part 1 of 2

ISS On-Orbit Status 2/24/03

All ISS systems continue to function nominally, except as noted 
previously or below. Day 93 in space for Expedition 6. Week 13 is 
underway.

The crew has half a day off in honor of Russia's "Defenders Day" 
holiday.

FE-1 Nikolai Budarin performed scheduled maintenance on the Elektron 
oxygen (O2) generator, replacing its BKO water purification column 
unit with a new spare. [The standalone BKO removes contaminants from 
the water to make it suitable for electrolysis into O2 and H2 
(hydrogen) by the Elektron. This BKO unit is identical with the BKO 
used in the condensate water recovery system (SRV-K). The purified 
water is stored in a dedicated EDV container (also referred to as the 
KOV) with a volume of 22 L. One liter of water produces about 25 
liters of O2, enough for one crewmember for one day.]

Budarin also turned off the gas analyzer in the Soyuz TMA-1 Descent 
Module which he had activated on 2/20 for air sampling.

CDR Ken Bowersox and FE-2/SO Don Pettit successfully completed the 
scheduled EVA DTO (development test objective), demonstrating the use 
of the U.S. EMU (extravehicular mobility unit) by two crewmembers 
without assist by a third IV person. The crew then tagged up with 
ground specialists to discuss results. [The three procedures performed 
by Sox and Don were EMU donning, in-suit prebreathe, and post-EVA. 
Suits were pressurized to verify ability to don SAFERs (simplified aid 
for EVA rescue) and exchange Metox (metal oxide) CO2 adsorbent 
canisters. After a successful ground demo on 2/14, some open questions 
had remained for the DTO to answer, such as: in case one of the EV 
crew is incapacitated after EVA, can EMU be doffed without any 
assistance? The demo showed it to be possible (albeit with some 
difficulty), by free-float "swimming" out of the HUT (hard upper 
torso). The DTO was accomplished inside the Airlock (A/L), with 
opening and closing of the IV hatch from the A/L Crewlock one of the 
demo objectives. To provide cooling for working in EMUs in the A/L, 
the ITCS LTL (internal thermal control system/low temperature loop) in 
the Node first had its RFCA's (rack flow control assembly's) flow 
setpoint raised (to 113.4 kg/hr); after the DTO, it was lowered again 
(45.4 kg/hr).]

Yesterday's replacement of the failed BVK-1 vacuum valve group on the
Vozdukh carbon dioxide (CO2) scrubber by Nikolai Budarin was 
successful.  Since yesterday afternoon, Vozdukh is up and running in 
Manual mode 5.

The U.S. CDRA (CO2 removal assembly) has been deactivated. [The check 
valve in adsorbent bed 2 apparently is stuck open, which required 
single-bed operation over the weekend until Vozdukh was recovered. 
Investigation continues.]

Two ppCO2 (CO2 partial pressure) readings were again taken by FE-2/SO 
Don Pettit with the CDMK (CO2 monitoring kit) in the Lab and SM, one 
in the morning, the other in the evening.

Bowersox completed the daily payload status inspection in the Lab, 
checking up on the PCG-STES010 (Protein Crystal Growth-Single Locker 
Thermal Enclosure System #10) payload.

All crewmembers performed their regular daily physical exercise (2.5 
hrs.) on TVIS treadmill, RED resistive device, CEVIS bike and VELO 
cycle plus load trainer.

R3 software upload operations in the U.S. segment have been 
successfully completed. Remaining activities are the corresponding 
upgrade of the Russian segment. Step-up of its onboard computer system 
(BVS) with the Version 7.01 software is scheduled for the week of 3/3, 
preceded by procedure review on 2/26.

Steps are being taken to correct an error detected in on-board 
emergency procedures for the case of a fire in the A/L, which 
automatically triggers a shutdown of the O2 system. The procedural 
change will assure that reactivation of the O2 system after the fire 
is extinguished will not reopen a hazardous situation.

Replacement of storage battery #1 in the FGB, which has been off-line 
for quite a while, is scheduled for tomorrow.

Today's targets for the CEO (crew earth observations program) were 
Taiwan Smog (this highly industrialized island generates a high volume 
of aerosols.  The ISS pass is ideal to view the entire island from the 
North to detect the development of smog plumes. Looking right of 
track), Lake Nasser, Toshka Lakes; Egypt (oblique context views of the 
new lakes growing to the west of the Nile are useful for monitoring 
change. Looking right of track), Seattle, Washington (where weather 
permits, the crew was to use this pass to document the snow pack of 
the Cascade Mountains to the right of track), and Denver, Colorado 
(although Denver itself may not have been visible this pass, crew was 
to look right of track to document the snow pack on the Front Range of 
the Rocky Mountains).

CEO images can be viewed at the website
http://eol.jsc.nasa.gov

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.